1. If a Ferrari, with an initial velocity of 10 m/s, accelerates at a rate of 50 m/s/s for 3 seconds, what will its final velocity be?

2.     While traveling along a highway a driver slows from 24 m/s to 15 m/s in 12 seconds. What is the automobile’s acceleration? (Remember that a negative value indicates a slowing down or

deceleration.)

3.     A cyclist accelerates at a rate of 7.0 m/s2. How long will it take the cyclist to reach a speed of 18 m/s?


1
Expert's answer
2021-11-30T09:49:36-0500

1)

"v=v_0+at,""v=10\ \dfrac{m}{s}+50\ \dfrac{m}{s^2}\times3\ s=160\ \dfrac{m}{s}."

2)

"a=\dfrac{v-v_0}{t}=\dfrac{15\ \dfrac{m}{s}-24\ \dfrac{m}{s}}{12\ s}=-0.75\ \dfrac{m}{s^2}."

3)

"v=v_0+at,""t=\dfrac{v-v_0}{a}=\dfrac{18\ \dfrac{m}{s^2}-0}{7\ \dfrac{m}{s^2}}=2.57\ s."